OpenStudy (anonymous):

Solve for x. 2 • 3 – 2 • 11 = 2(x – 11) I'm doing this one by Order of operations, at least I think I should, There are four answer choices, do I just plug them in and see if its correct?

OpenStudy (anonymous):

you can plug if there multiple choice question... but... 6-22=2x-22 -> 6=2x -> x=3

OpenStudy (anonymous):

Yay, because it is multiple choice, Thanks :) @purger

OpenStudy (anonymous):

2 * 3 - 2 * 11 = 2(x - 11) 6 - 22 = 2x - 22 - 16 = 2x - 22 - 2x = - 6 x = 3

OpenStudy (anonymous):

Thank you too! ^-^ @britneyspears

OpenStudy (anonymous):

You're welcome!! Xo!